Find each of the following percentile points and draw the appropriate normal



curve. Complete your procedures.



1.The results of the entrance examination for freshmen are normally distributed with x=87 and s=12.5. What is the percentile rank of a score of 90?

1
Expert's answer
2022-04-10T17:03:03-0400

"P(X<90)=P(Z<\frac{90-87}{12.5})=P(Z<0.24)=0.59483"